# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ 10       $tentative\coloneqq g(\mathcal{G}_{\text{current}})+1$;
12          $g(\mathcal{G}_{\text{neighbor}})\coloneqq tentative$;
13            add $\mathcal{G}_{\text{neighbor}}$ to $openSet$ with priority $g(\mathcal{G}_{\text{neighbor}})+h(\mathcal{G}_{\text{neighbor}})$;
Algorithm 3 An A* algorithm for computing the up-down distance $d_{\L ,\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})$
Finding a lower-bound for $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$: We next present the algorithm for computing the lower bound for $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$ according to Property 2. Algorithm 4 presents the algorithm which uses Algorithm 5.
input : CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ among nodes $V$, functions getDisjointV and findMinOps
output : The lower bound for the $\mathcal{M}$-oriented distance between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$
1 begin
2 $\mathcal{G}_{t,\text{undir}}\coloneqq undirect(\mathcal{G}_{t})$
3 $\mathcal{G}_{s,\text{undir}}\coloneqq undirect(\mathcal{G}_{s})$
 $d\coloneq|\mathrm{sk}(\mathcal{G}_{t,\text{undir}})\triangle\mathrm{sk}(\mathcal{G}_{s,\text{undir}})|$ ;
 // skeleton difference
 // Get all v structures
 $vList\coloneqq[]$ ;
 // list of uncommon undirected v-structures in the skeletons of the two graphs
4 foreach node $i\in V$ do
5    foreach node $j\in V$ do
6       foreach node $k\in V$ do
7          if $i\neq j\land j\neq k\land i\neq k\land((i,j)\in E(\mathcal{G}_{t,\text{undir}})\land(j,k)\in E(\mathcal{G}_{t,\text{undir}}))\land((i,j)\in E(\mathcal{G}_{s,\text{undir}})\land(j,k)\in E(\mathcal{G}_{s,\text{undir}}))\land(i,k)\notin E(\mathcal{G}_{t,\text{undir}})\land(i,k)\notin E(\mathcal{G}_{s,\text{undir}})$ then
8             if $((i,j)\in E(\mathcal{G}_{t})\land(k,j)\in E(\mathcal{G}_{t})\land(j,i)\notin E(\mathcal{G}_{t})\land(j,k)\notin E(\mathcal{G}_{t})\land((i,j)\notin E(\mathcal{G}_{s}))\lor(k,j)\notin E(\mathcal{G}_{s})))\lor((i,j)\in E(\mathcal{G}_{s})\land(k,j)\in E(\mathcal{G}_{s})\land(j,i)\notin E(\mathcal{G}_{s})\land(j,k)\notin E(\mathcal{G}_{s})\land((i,j)\notin E(\mathcal{G}_{t}))\lor(k,j)\notin E(\mathcal{G}_{t})))$ then
9                  append v structure $(i,j,k)$ to vListvList
 $groups\coloneqq getDistjointVGroups(vList)$ // get v structure groups with no shared vertices
10 $h\coloneqq 0$
11 if $|vList|>0$ then
12    foreach $group\in groups$ do
13       $h\coloneq h+2(findMinOps(group))$
14 return h + d
Algorithm 4 Algorithm for computing the lower bound for the $\mathcal{M}$-oriented distance according to Property 2.
input : Set $S$ consisting of a collection of triplets of vertices, operations $\mathrm{OP}$
output : Returns the minimum number of operations to remove all elements in $S$
1 begin
 $queue\coloneqq[(S,0)]$ // tracks state and step count
 $visited\coloneqq\{\}$ // avoid repeat states
@@ 13       $h\coloneq h+2(findMinOps(group))$
3 while $queue\neq\emptyset$ do
4    $current\coloneqq$ get first element of the queue
5      remove first element of queuequeue
6
7    $subS\coloneqq$ subset of S in currentcurrent
8    $steps\coloneqq$ steps in currentcurrent
9    if $subS$ is empty then
10       return steps
    // Get all pairs (i, j) from triples in S
    $pairs\coloneqq{}$ // set of unique pairs from triples
11    foreach $triple\in subS$ do
12         add all combinations of 2 vertices from the tripletriple of vertices to pairspairs
13    foreach $pair\in Lairs$ do
14       $i\coloneqq$ first node in pairpair
15       $j\coloneqq$ second node in pairpair
@@
17       foreach $op\in\mathrm{OP}$ do
18          $newS\coloneqq op(subS,i,j)$
19          if $newS\notin visited$ then
20               add $newS$ to visitedvisited
21               add $(newS,steps+1)$ to queuequeue
 return NULL // No Solution
22
Algorithm 5 Algorithm findMinOps used in Algorithm 4.

### v2

@@ 12             $\textnormal{{Push}}(\mathrm{S}^{\ast},(a,b,c))$
18 foreach $\mathrm{S}\in\textnormal{{DisjointTriplets}}(\mathrm{S}^{\ast})$ do
19 $h\leftarrow h+2\cdot\textnormal{{FindMinOps}}(\mathrm{S})$
return $h+d$
Algorithm 2 Computing the lower bound in Proposition˜6 on $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})$.
Input: Set $\mathrm{S}$ consisting of a collection of vertex triplets
Output: $m_{\text{op}}(\mathrm{S})$
$\textnormal{{Q}}\leftarrow\textnormal{{Dict}}()$
// tracks state and step count
1 $\textnormal{{Q}}[S]\leftarrow 0$
@@
3while Q is not empty do
4 $(S^{\prime},k)\leftarrow\textnormal{{Pop}}(Q)$
5 if $S^{\prime}=\emptyset$ then
6    return $k$
7 $\textnormal{{pairs}}\leftarrow\{\}$
8 foreach $(a,b,c)\in S^{\prime}$ do
9    $\textnormal{{Push}}(\textnormal{{pairs}},(a,b))$; $\textnormal{{Push}}(\textnormal{{pairs}},(a,c))$; $\textnormal{{Push}}(\textnormal{{pairs}},(b,c))$
10
11 foreach $(i,j)\in\textnormal{{pairs}}$ do
12    foreach $\mathrm{op}\in\left\{\mathrm{OP}_{(i,j),\mathrm{CE}},\mathrm{OP}_{(i,j),\mathrm{ER}}\right\}$ do
13       $S^{\prime\prime}\leftarrow\mathrm{op}(S^{\prime})$
14       $\textnormal{{Q}}[S^{\prime\prime}]\leftarrow k+1$
15
@@ 14       $\textnormal{{Q}}[S^{\prime\prime}]\leftarrow k+1$
return 1-1
// No Solution
Algorithm 3 Subroutine FindMinOps for computing $m_{\text{op}}(\cdot)$ in Eq.˜6.
I.3 Additional numerical results
Fig.˜S14 compares the $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$ and the structural Hamming distances for polytree MPDAGs over 5 vertices.
Figure S14: The joint distribution of $d_{\mathcal{L},\operatorname{\mathrm{pseudo}}}$ (7) and the structural Hamming distance computed for all pairs of polytree MPDAGs over 5 vertices, where the identity line is drawn as dashed.